9x-9y=27
7y-3x=-11
Solve by the elimination method..
Solution?

Answer by scott8148(6628)

dividing 1st eqn by 3 ___ 3x - 3y = 9

adding eqns to ELIMINATE x ___ 7y - 3x + 3x - 3y = -11 + 9

4y = -2 ___ y = -1/2

substituting ___ 3x - 3(-1/2) = 9 ___ x = 5/2